OpenStudy (anonymous):

Change to polar coordinates: 2y=x

OpenStudy (phi):

use x= r cos A and y = r sin A

OpenStudy (anonymous):

so is it 2rsinA = rcosA @phi

OpenStudy (phi):

yes but you can simplify that to tan A = 1/2

OpenStudy (anonymous):

@phi why? i dont understand how it simplifies to tan A = 1/2

OpenStudy (phi):

2 r sin A = r cos A divide both sides by r 2 sin A = cos A divide both sides by cos A 2 (sin A)/(cos A)= 1 divide both sides by 2 sin A/ cos A= 1/2 also, tan A = sin A/ cos A

OpenStudy (phi):

we can write this as \[ A = \tan^{-1} \frac{1}{2} \]

OpenStudy (anonymous):

@phi alright i get it, thanks!